OpenStudy (anonymous):

use the ratio 20ft/1 ft = 8.5heads/x heads to find the hand size of the new human

OpenStudy (anonymous):

can you help me?

OpenStudy (dumbcow):

same type of problem cross multiply

OpenStudy (anonymous):

is but can you solve it

OpenStudy (dumbcow):

yes but i want you to be able to solve it anytime you have 2 fractions equal to each other multiply each numerator by the other denominator and set the 2 products equal to each other

OpenStudy (dumbcow):

Given A/B = C/x A*x = B*C x = (B*C)/A in this case A=20 B=1 C=8.5

OpenStudy (anonymous):

20/1=8.5/1 x=0.425